Find the mean and variance of the random variable X with probability function or density f(x)

f(x) = k(1 – x2) if –1 3x = 1 and 0 otherwise
Answer #1

The value of k is got by noting that the Total Probability = 1.

Thus, we get:

\int_{-1}^{1}k(1-x^{2})dx=1

k(x-\frac{x^{3}}{3})=1

between limits -1 to 1.

Applying limits, we get:

\frac{4k}{3}=1

So,

k = 0.75

So,

The probability density function is:

f(3) = 0.75(1-2),

                           for -1\leqx\leq1

Mean = E(X) is got as follows:

E(x)=\int_{-1}^{1}0.75x(1-x^{2})dx

=0.75(\frac{x^{2}}{2}-\frac{x^{4}}{4})

between the limits - 1 to 1

Applying limits, we get:
E(X) = 0

E(X2) is got as follows:

E(x^{2})=\int_{-1}^{1}0.75x^{2}(1-x^{2})dx

=0.75(\frac{x^{3}}{3}-\frac{x^{5}}{5})

Applying limits, we get:

=0.75(\frac{2}{3}-\frac{2}{5})=0.2

Variance = E(X2) - (E(X))2

                = 0.2 - 02

              = 0.2

So,

Answers are:
Mean = 0

Variance = 0.2